OpenStudy (anonymous):

1. 4 ⋅ 6 + 5 ⋅ 7 + 6 ⋅ 8 + ... + 4n( 4n + 2) = The "n's" are supposed to be lower than the number

OpenStudy (anonymous):

Is this asking for a sum? \[\Large \sum_{n=1}^N 4n(4n+2)= \]

OpenStudy (anonymous):

seems very hard to understand where this question is going, and furthermore \(4_n(4_n+1)\) , does not yet make much sense, except you have some recursive formula and then it should most likely be \(a_n\)? but I might digress from the question itself.

OpenStudy (loser66):

@spacelimbus I don't see 4n(4n+2) is general form for the terms in sequence

OpenStudy (anonymous):

probably this: \[\Large \sum_{n=4}^{4n} n(n+2)=\sum_{n=4}^{4n}(n^2+2n) =\sum_{n=4}^{4n}n^2+2\sum_{n=4}^{4n}n\]

OpenStudy (loser66):

except the first term, 4*6,

OpenStudy (anonymous):

Yes @Loser66 , I was paying too much attention to the last term of the given sum, however the above should fix this :) Only requires a shift of index and plugging in.

OpenStudy (loser66):

I think we have to wait for him to make it clear.

OpenStudy (anonymous):

Use mathematical induction to prove the statement is true for all positive integers n, or show why it is false.

OpenStudy (anonymous):

sorry for the wait on me. i had to run out real quick

OpenStudy (anonymous):

but there is no statement yet is there?

OpenStudy (anonymous):

You should check and update the precise question again @seesawn, see the thing you have in your opening post is nothing but a mathematical expression, to use induction you would to have at least one closed form of this expression to which it is equal to. If you don't have this, you can compute one on your own using summation. However, it would be a pointless task to use the summation method and then prove it again with induction, so I assume that you're question is missing something.

OpenStudy (anonymous):

Use mathematical induction to prove the statement is true for all positive integers n, or show why it is false. 1. 4 ⋅ 6 + 5 ⋅ 7 + 6 ⋅ 8 + ... + 4n( 4n + 2) = thats my question

OpenStudy (anonymous):

I am afraid I have to leave this open then, since this is not a question, you can see that for yourself that there is no right hand side next to the equal sign, which in fact makes it obsolete, they could have just written \(4 \cdot 6 + 5 \cdot 7 + 6 \cdot 8 + \dots + 4n(4n+2)\) as well with the same information.

OpenStudy (anonymous):

the equation is supposed to look like this \[4_{n}(4_{n}+2)\]

OpenStudy (anonymous):

Good, I am bereft of understanding this question :) I suggest to you to bump it to the top once in a while on OpenStudy and see if someone answers that understands it some more. Best luck!

OpenStudy (anonymous):

thank you for trying!!!

OpenStudy (anonymous):

your sequence should be something like this: \[4\cdot 6 + 5\cdot 7 + 6\cdot 8 + \cdots + (4+n)(6 + n) \text{ where n starts at }0\] i'm not sure what your \(4_n\) is supposed to mean.

OpenStudy (anonymous):

im just showing it how i was shown it on the paper

OpenStudy (anonymous):

as such, you could rewrite the terms as \(24 + 10n +n^2,n\in \mathbb{Z} \). then the sum you could work from your formulas. \[\sum_{i=0}^{n}\left(i^2+10i+24\right)=\sum_{i=0}^{n}i^2+\sum_{i=0}^{n}10i+\sum_{i=0}^{n}24=\frac{n(n+1)(2n+1)}{6}+\frac{n(n+1)}{2}+24(n+1)\]

OpenStudy (anonymous):

the last term being 24(n+1). this is for n starting at 0.

OpenStudy (anonymous):

i understand that what is shown on the paper, but it's quite useless unless you undersatnd what it means. I don't know what it means so it doesn't do me any good.

OpenStudy (anonymous):

my math homework. i am just as confused as everyone that is trying to help me

OpenStudy (anonymous):

no it isnt. its from pre calculus. making up credits online

OpenStudy (anonymous):

yes way. im a senior in high school, screwed off my sophmore year and have to make up the credit. im freaking confused

OpenStudy (anonymous):

it's (4 times 6) plus (5 times 7) plus ..., yes?

OpenStudy (anonymous):

idk, im going to try that and see if it works

OpenStudy (dan815):

hey

OpenStudy (dan815):

I can solve your question for ya

OpenStudy (dan815):

well you see if i were to write a couple of sums

OpenStudy (dan815):

and take differences of hte sums

OpenStudy (dan815):

the 3rd difference is constant and is =32

OpenStudy (anonymous):

?

OpenStudy (dan815):

this means your solution is in the form sum(x)=ax^3+bx^2+cx+d

OpenStudy (dan815):

you can just plug in for 4 different independant equations and solve your unknowns

OpenStudy (dan815):

or you can simplify it starting with the 32

OpenStudy (anonymous):

24, 59, 107, 170, ... where are the differences the same?

OpenStudy (anonymous):

the difference is the next term

OpenStudy (dan815):